Main INDEX
Monthly INDEX
PREV
NEXT
User name Feuerbach
Log entry time 17:02:43 on May 12, 2004
Entry number 121501
keyword=May 12 Shift Plan
shiftplan
E94-107 Shift Plan: Wednesday, May 12, 8:00 am -- Thursday, May 13, 4:00 pm
May 12, 4:00 pm, Robert Feuerbach, Run Coordinator (pager 584-5411)
Continue taking data in the current kinematic setting on Be-9:
- Beam: E=3.7 (MBSY1C Bdl=943721; check after trips); up to 120uA; FB ON; raster ON at 0.5 x 0.5 mm^2 (full width from spot++); page RC if SLI energy spread >= 3.E-5
- Target: thin 9Be (this is the "thick" target on the target controls)
- Left HRS: P0=1.96 GeV/c; Iseptum=181.46 A (this is magnet current, not PS; monitor it)
- Right HRS: P0=1.57 GeV/c; Iseptum=145.35 A
Scheduled events/Special tasks:
- The beam position at the dump-viewer has jumped to the far edge, and might need to be recentered upon restoration.
- During the Wednesday, Thursday, and Friday (May 12, 13, 14) DAY shifts, Hall C will conduct leakage tests. These shifts are on STANDBY. If the accelerator is up but Hall C cannot take data, the shift will become ACTIVE: in this case please contact the Run Coordinator.
General Tasks:
- When recording settings, check whether they are correct; don't just copy them.
- For each run do detector replay and compare with reference plots.
- After end of run do full physics replay from the ana directory.
- HALOG any detector, DAQ trip and unusual observation when they happen.
- Do shift checklist.
- Twice a shift HALOG control screens and septum & energy spread strip charts.
- For STANDBY shifts, please HALOG the control screens and magnet strip charts.
- If septa Tcoil >= 7.5 K, HALOG with TECH in title; only 8:00 am - 5:00 pm also call Ed Folts.
- Any problem that prevents data taking and that you cannot solve within 30 min page the Run-Coordinator.
- During night and weekend shifts contact Tech on-call only for emergencies (in doubt page RC). HALOG all events that will require TECH assistance with the word "TECH" in the title.
Known Problems
- N2 level reading right Q3 (CLL72Q310) is not working properly; ignore
- Dump diffuser cooling water circuit might trigger FSD; after hours page Run Coordinator
Please log below (and in the HALOG) any unusual observations and problems solved
Date/Time |
Name |
Comment |
HALOG |
|
|
|
|